A light-rail commuter train accelerates at a rate of 1.35 m/s2. How long (in s) does it take to reach its top speed of 80.0 km/h, starting from rest?
1
Expert's answer
2020-02-10T09:34:41-0500

The acceleration by definition


"a=\frac{\Delta v}{\Delta t}."

Hence


"\Delta t=\frac{\Delta v}{a}=\frac{80.0\times \frac{1000\rm \: m}{3600\:\rm s}}{1.35\:\rm m/s^2}=16.5\:\rm s."